OpenStudy (anonymous):

the volume of a cylindrical can is 200 cubic centimeters. its surface area S as a function of the radius r can be given by S(r)=2(pi)r^2 + 400/r. Write S over a common denominator and find the S(4)

jhonyy9 (jhonyy9):

so than v=200 cm^3 from this result that 200 = the height time area of base so 200=h*pir^2 --- from this result that pir^2 =200/h - so if S(r)=2pir^2 +400/r =2(200/h) +400/r =400/h +400/r =400(r+h)/rh - so S(r)=400(r+h)/rh S(4)=400(4+h)/4h =100(4+h)/h